Obtain the complete integral xp^2- ypq + y^3q - y^2z= 0


1
Expert's answer
2021-01-18T15:52:29-0500

Using charpit's auxilliary equation,


"\frac{dp}{∂f/∂x + p(∂f/∂z)}=\frac{dq}{∂f/∂y + q(∂f/∂z)}=\frac{dz}{−p(∂f/∂p)− q(∂f/∂q)}=\frac{dx}{−∂f/∂p}=\frac{dy}{−∂f/∂q}\\

\frac{∂f}{∂x}=p^2\\
\frac{∂f}{∂z}=-y^{2}\\
\frac{∂f}{∂y}=-pq+3qy^2-2zy\\
\frac{∂f}{∂p}=2xp-yq\\
\frac{∂f}{∂q}=-yp+y^3\\

\text{so by substitution}\\

\frac{dp}{p^2-py^2}=\frac{dq}{2qy^2-pq-2zy}=\frac{dz}{-2xp^2+2ypq-qy^3}=\frac{dx}{yq-2xp}=\frac{dy}{yp-y^3}\\

\frac{dp}{p}-\frac{dy}{y}=0\\
log p-logy=log a\\
log p=log a+logy\\
p=ay\\
\text{substitute the value of p in the equation and we have q to be}\\

x({ay})^2-y(ay)q+q({ay})^3-zy^2=0\\
q=\frac{z-xa^2}{ay-a}\\
\text{So we use}\\ 
dz=pdx+qdy\\
dz=aydx+ \frac{z-xa^2}{ay-a}dy\\
\text{by integrating both sides, we have}\\
z=axy+(z-xa^2)log_e(ay-a)\\
z=axy+blog_e(ay-a)\\
\text{Where a and b are arbitrary constants}"